# Observational constraints on degeneracy of non-canonical inflation driven by quartic potential

### v2

@@ V Conclusion
In this work, we investigated the quartic potential in non-canonical inflationary model with power-law Lagrangian (2). We showed that, in contrary to the standard inflation, the predictions of this potential in non-canonical framework are consistent with the observational data of Planck 2018. Among the free parameters of the model, only the changes of $\alpha$ parameter have a significant effect on the predictions of the model.
We checked the model based on the changes of $\alpha$ considering different values for the duration of the inflation epoch $N$. To determine the allowed range of $\alpha$ parameter, we used theoretical considerations as well as observational constraints. We showed that the sound speed parameter $c_{s}$, equilateral non-Gaussianity $f_{\rm NL}^{\rm equil}$, scalar spectral index $n_{s}$, and tensor-to-scalar ratio $r$ bound the parameter $\alpha$ (see Table 1). The minimum value of $\alpha$ was determined to be larger than one based on the constraint on $c_{s}$. The equilateral non-Gaussianity constraint also limits the maximum value of $\alpha$ to be $\alpha\leq 130$. The $r-n_{s}$ curve sets the lower bound on $\alpha$ considering specific duration for inflation. For example, for $N=60$ we obtain $\alpha\geq 18.8$.
Then, we examined the observational predictions of the model for $n_{s}$ and $r$ with respect to $\alpha$ parameter. It was found that, the scalar spectral index of this model shows degeneracy at large values of $\alpha$ parameter (see Fig. 1). In the following, we examined three parameters of the reheating epoch, including the equation of state parameter $\omega_{re}$, the duration of the reheating epoch $N_{re}$, and the temperature of the reheating epoch $T_{re}$. We examined $N_{re}$ and $T_{re}$ in terms of $n_{s}$ for several values of $\alpha$. We showed that by varying $\alpha$, the obtained values for both $N_{re}$ and $T_{re}$ in terms of $n_{s}$ do not exhibit much differentiation and are somewhat degenerate. The condition $N_{re}\geq 0$, constrains the length of the inflation epoch as $N\leq 55.7$. On the other side, from the $r-n_{s}$ diagram we have $N\geq 55$. Altogether, the condition $N_{re}\geq 0$ thereto $r-n_{s}$ observational data specified the allowed length of the inflationary period as $55\leq N\leq 55.7$ and subsequently the non-canonical parameters $\alpha$ and $\mu$ were constrained (see Table 2). For $N=55$, we have $110\leq\alpha\leq 130$ and for $N=55.7$, $60\leq\alpha\leq 65$ is obtained.
In the following, we examined the relic GWs for this model. We checked the present density of the relic GWs for different values of $\alpha$ parameter and duration of inflation between 55 to 55.7 $e$-folds. We found that, the present density parameter of the relic GWs is sensitive to changes in $\alpha$ parameter. So using the relic GWs considerations, the degeneracy of the model with respect to $\alpha$ parameter can be broken. Moreover, we showed that the present density of relic GWs for this model, falls within the sensitivity ranges of developing GWs detectors.

### v3

@@ V Conclusion
In this work, we investigated the quartic potential in non-canonical inflationary model with power-law Lagrangian (2). We showed that, in contrary to the standard inflation, the predictions of this potential in non-canonical framework are consistent with the observational data of Planck and BICEP/ Keck 2018. Among the free parameters of the model, only the changes of $\alpha$ parameter have a significant effect on the predictions of the model.
We checked the model based on the changes of $\alpha$ considering different values for the duration of the inflation epoch $N$. To determine the allowed range of $\alpha$ parameter, we used theoretical considerations as well as observational constraints. We showed that the sound speed parameter $c_{s}$, equilateral non-Gaussianity $f_{\rm NL}^{\rm equil}$, scalar spectral index $n_{\rm s}$, and tensor-to-scalar ratio $r$ bound the parameter $\alpha$ (see Table 1). The minimum value of $\alpha$ was determined to be larger than one based on the constraint on $c_{s}$. The equilateral non-Gaussianity constraint also limits the maximum value of $\alpha$ to $\alpha\leq 130$. The $r-n_{\rm s}$ curve sets the lower bound on $\alpha$ considering specific duration for inflation. For example, for $N=60$ we obtain $\alpha\geq 18.8$.
Then, we examined the observational predictions of the model for $n_{\rm s}$ and $r$ with respect to $\alpha$ parameter. It was found that, the scalar spectral index of this model shows degeneracy at large values of $\alpha$ parameter (see Fig. 2). In the following, we examined three parameters of the reheating epoch, including the equation of state parameter $\omega_{\rm re}$, the duration of the reheating epoch $N_{\rm re}$, and the temperature of the reheating epoch $T_{\rm re}$. We showed that the degeneracy of the model with respect to the parameter $\alpha$ is broken by analyzing $N_{\rm re}$ and $T_{\rm re}$ as functions of $\alpha$ for fixed values of the $e$-folds number $N$. The condition $N_{\rm re}\geq 0$ impose an upper bound on the length of the inflation epoch as $N\leq 55.7$. On the other side, from the $r-n_{\rm s}$ diagram we have $N\geq 55$. Altogether, the condition $N_{\rm re}\geq 0$ thereto $r-n_{\rm s}$ observational data specified the allowed length of the inflationary period as $55\leq N\leq 55.7$ and subsequently the non-canonical parameters $\alpha$ and $\mu$ were constrained (see Table 1). For $N=55$ and $N=55.7$, the allowed ranges of the $\alpha$ parameter are obtained as $110\leq\alpha\leq 130$ and $60\leq\alpha\leq 65$, respectively.
In addition, we examined the relic GWs for this model. We checked the present density of the relic GWs for different values of $\alpha$ parameter and duration of inflation between 55 to 55.7 $e$-folds. We showed that the present density of relic GWs for this model, falls within the sensitivity ranges of developing GWs detectors.
@@ Acknowledgements
Data Availability Statement
No Data associated in the manuscript.